### v2

# SCILLA: SurfaCe Implicit Learning for Large Urban Area, a volumetric hybrid solution

Abstract
Neural implicit surface representation methods have recently shown impressive 3D reconstruction results. However, existing solutions struggle to reconstruct urban outdoor scenes due to their large, unbounded, and highly detailed nature. Hence, to achieve accurate reconstructions, additional supervision data such as LiDAR, strong geometric priors, and long training times are required. To tackle such issues, we present SCILLA, a new hybrid implicit surface learning method to reconstruct large driving scenes from 2D images. SCILLA’s hybrid architecture models two separate implicit fields: one for the volumetric density and another for the signed distance to the surface. To accurately represent urban outdoor scenarios, we introduce a novel volume-rendering strategy that relies on self-supervised probabilistic density estimation to sample points near the surface and transition progressively from volumetric to surface representation. Our solution permits a proper and fast initialization of the signed distance field without relying on any geometric prior on the scene, compared to concurrent methods. By conducting extensive experiments on four outdoor driving datasets, we show that SCILLA can learn an accurate and detailed 3D surface scene representation in various urban scenarios while being two times faster to train compared to previous state-of-the-art solutions.
Keywords:
Neural Implicit Surface Reconstruction Radiance Fields Mesh Reconstruction
[…]
Figure 1: We introduce SCILLA, a novel hybrid implicit learning method capable to accurately reconstruct large driving scenes from 2D input images solely. Extensive experiments on four popular driving datasets show the superiority of SCILLA’s mesh (left) over previous state-of-the-art methods such as StreetSurf[11] (right). Figure shows both original and textured mesh obtained from both methods.
1 Introduction
Achieving accurate 3D reconstruction from multi-view-consistent images has been an important topic since decades in computer vision and computer graphics communities. A 3D representation of the perceived environment is fundamental for addressing many tasks, such as scene re-lighting [41], scene editing [42] and 3D object insertion for artificial data augmentation [36].
Recently, neural implicit surface representation methods[34, 39] have shown impressive 3D surface reconstruction performance from monocular images, achieving highly detailed and precise mesh representations. Such methods are originally inspired by Neural Radiance Fields (NeRF) [21], and reframed for multi-view surface reconstruction. NeRF introduced volume rendering for Novel View Synthesis (NVS) by sampling points along camera rays. These points, along with their viewing directions, are fed into a Multi-Layer Perceptron (MLP) to predict their respective colors and densities. The final image is rendered by blending these predicted values using alpha compositing. Conversely, neural surface representation methods aim to approximate the surface using a learnable Signed Distance Function (SDF) [34, 39]. Unlike NeRF, the outputs of the MLP are colors and SDF values instead of colors and volumetric densities, which are later blended using a modified alpha-compositing equation adapted to SDF. An accurately reconstructed surface is represented as a zero-level set of the SDF. Assuming the objects lying at the center of the scene, the SDF is first initialized to form a spherical shape to bound the region of interest. Although current methods achieve impressive reconstruction results, they are mostly applied to simple, bounded and object-centric scenes. Hence, applying such methods to unbounded urban scenarios results in sub-optimal reconstruction given their significantly different nature. To overcome such limitations, implicit surface reconstruction approaches commonly rely on LiDAR supervision [25, 36] or strong geometric priors [11] and require longer training times to produce qualitative outdoor scene reconstruction meshes from 2D images.
To address those challenges, we propose SCILLA, a hybrid approach to accurately initialize a signed distance field of an unbounded area and represent complex surfaces from urban outdoor scenarios with high accuracy and geometric precision while requiring half of the training time compared to other state-of-the-art solutions such as StreetSurf [11]. We propose a hybrid architecture that models two separate implicit fields: one representing the volumetric density of the scene and another one representing the signed distance to the surface. SCILLA’s novel volume rendering approach relies on a self-supervised density estimator which permits to sample points near the surface along with a progressive sampling strategy that smoothly transitions from a volumetric to a surface representation, drastically reducing the model convergence time. Inspired by previous state-of-the-art works [11, 40], we supervise surface normals by relying on normal cues predicted from a pre-trained monocular estimator [8]. We extensively validate our solution on KITTI-360 [18], Pandaset [38], Waymo Open Dataset [29], and nuScenes [3]. By comparing the quality of the extracted meshes from the signed distance field, we show that our method can efficiently reconstruct urban outdoor scenarios showing SoTA accuracy. To summarize, the main contributions of our work are:
  • A novel implicit hybrid learning method for surface reconstruction on large outdoor driving scenes,
  • A hybrid architecture that models two distinct implicit fields to represent both the volumetric density and the signed distance field,
  • A new volume rendering technique that relies on self-supervised probabilistic density estimation to efficiently sample points near the surface and smoothly transition from a volumetric representation to a surface representation, significantly reducing training times,
  • Various regularization techniques of the signed distance field to enhance our model’s implicit surface representation learning and convergence are presented.
2 Related work
Multi-view surface reconstruction.
Traditional surface reconstruction techniques represent a scene either by occupancy grids [7, 24, 31, 32] or relying on point clouds [9, 27, 28, 26]. Occupancy grids-based methods are often limited by the voxel resolution and suffer from high memory consumption. On the other hand, Multi-View Stereo (MVS) techniques [22] generate, from posed images, dense point clouds associated with surface normal that can be utilized in a subsequent surface reconstruction step [14]. These methods heavily depend on the quality of the generated point cloud. Nevertheless, they can accumulate errors due to their complex multi-step pipelines, and are mostly designed for object-centric scenes with high overlaps between the observations. We refer the reader to dedicated surveys for in-depth analysis of such techniques [43, 20, 2].
Neural implicit surface representations.
NeRF-based methods can achieve impressive NVS results but fail to accurately reconstruct scene surfaces given the lack of constraints or priors on the surface when modelling the scene as a density volume. Neural surface reconstruction approaches on the other hand, explicitly represent the surface as an SDF and propose various adaptations to train such methods through differentiable volume rendering [34, 39]. NeuS [34] proposes a novel SDF-based rendering equation that relies on SDF values to compute the alpha blending factor used in volume rendering instead of densities as presented in the original NeRF [21]. Nevertheless, NeuS’s original contribution is designed mainly for object-centric datasets and takes long training times. Some recent works aim to reduce training times [35] or reconstruct large-scale highly detailed landmark scenes [17]. However, such approaches are not tailored for large unbounded driving scenarios and still require relatively long training times. In contrast, our method, is built upon NeuS’s formulation and it is designed to reliably represent urban outdoor scenes.
Urban outdoor surface reconstruction.
Urban outdoor scene reconstruction is one of the main subjects contributing to resolving many downstream applications, such as inverse rendering and object insertion [36]. Latest state-of-the-art approaches rely on SDF-based surface reconstruction for urban outdoor scenes, to disentangle the scene’s geometry from other scene parameters.
Existing SoTA methods relying on neural implicit surface reconstruction are mainly guided by LiDAR [36, 25], limiting their large-scale deployment.
One of the few existing approaches that tackle the subject of surface reconstruction of urban outdoor scenes is StreetSurf[11]. The authors of StreetSurf present a solution that disentangles close-range view from distant-range view, alongside a signed distance field initialization technique to match the road surface according to the ego poses of the vehicle. This surface initialization approach is rough, not straightforward, and heavily depends on the ego poses and the assumption that the road and the sidewalk are at the same height in every part of the scene. Additionally, StreetSurf relies on the assumption that self-driving scenes are mostly long and narrow and showcase reconstruction of the scene from only one public dataset (i.e. Waymo Open Dataset [29]). In reality, driving datasets frequently contain scenarios where the scenes are short and curvy, hence, disentangling the close and far range in those situations becomes more challenging. Instead, our solution is more adaptive in comparison to StreetSurf as it does not rely on strong assumptions about the scene’s topology. Furthermore, SCILLA is also more compact and faster to train as the entire scene is represented by a single hybrid implicit field.
Hybrid scene representation.
Some state-of-the-art methods have already investigated the use of hybrid rendering solutions relying on both volumetric and surface composing. Wang et al. [37] introduced a hybrid rendering method based on NeuS [34] that uses a spatially varying kernel, which changes according to the fuzziness of the surface. Based on the kernel, a narrow mesh envelope is extracted around the surface, which then enables ray casting only inside this envelope. This solution helps accelerate both rendering time and quality. Other methods [12, 13, 30] focus on a hybrid representation that both relies on an explicit representation (meshes) and an SDF representation. The hybrid volume-mesh representation enables efficient rendering and less resourceful storage while outputting an implicit geometrical representation. In addition to fast training times, this representation also benefits from easy integration into computer graphics engines. NeuSG [5] combines 3D Gaussian Splatting [15] with NeuS to obtain an accurate Gaussian cloud close to the surface that serves as a direct supervision signal during NeuS training. The aforementioned solutions mainly focus on improving the rendering speed and are better fitted for small scenes with a hybrid representation. In our proposal, we jointly train a hybrid model to benefit from the fast training of a density-based approach along with an accurate geometric representation obtained thanks to the signed distance field learning.
3 Method
SCILLA models the volumetric density and signed distance field in a hybrid architecture. The details of this architecture are presented in Sec. 3.2, while the novel underlying volume rendering strategy is explained in Sec. 3.3. Our hybrid strategy relies on a self-supervised density distribution estimation to sample points near the surface and progressively transition from a volumetric to an implicit surface representation. To facilitate the convergence of our method, we further introduce regularization techniques for the signed distance field. The overall architecture of our proposed method is illustrated in Fig. 2.
[…]
Figure 2: SCILLA overview – Our solution can be divided into two key components: an hybrid scene representation (Sec. 3.2) and a ray-based volumetric rendering that progressively transitions from density to SDF sample alpha composition (Sec. 3.3).
3.1 Preliminaries
3.1.1 NeRF’s Implicit Volumetric Representation.
NeRF [21] takes as input the 3D positions $x_{i}\in\mathbb{R}^{3}$ and the viewing direction $d\in\mathbb{R}^{2}$ of sampled points along each camera ray $r$ and predicts their corresponding densities $\sigma_{i}\in\mathbb{R}$ and colors $c_{i}\in\mathbb{R}^{3}$. This representation of the scene is approximated using a Multi-Layer Perceptron (MLP). Finally, volume rendering is used to alpha-composite the colors of the samples along each ray and render the final pixel color $\hat{C}(r)\in\mathbb{R}^{3}$:
$\hat{C}(r)=\sum_{i=1}^{N}w_{i}c_{i},\ \ w_{i}=T_{i}\alpha_{i},$ (1)
where $T_{i}=\prod_{j=1}^{i-1}(1-\alpha_{j})$ is the accumulated transmittance along the ray and $\alpha_{i}\in\mathbb{R}$ a blending factor. This factor (referred to in the literature as the alpha-value) is used for alpha-composing, and is given by:
$\alpha_{i}=1-\exp(-\sigma_{i}\delta_{i}),$ (2)
with $\delta_{i}\in\mathbb{R}$ being the distance between the samples along the ray. Note that in SCILLA’s hybrid formulation, we refer to this value as $\alpha^{v}_{i}$ to disambiguate from alpha values derived from the estimated SDF which is introduced in Eq. 3.
3.1.2 SDF representation: NeuS.
For an accurate surface-based representation of a 3D scene, NeuS [34] presented a different formulation of the NeRF’s volume rendering equation. They represent the scene as a signed distance function, where the surface is the zero-level set of such function: $S=\left\{x\in\mathbb{R}^{3}|f(x)=0\right\}$, where $f(x)$ is the signed distance value at a given position $x$. To efficiently represent surfaces, NeuS adopted a volume rendering formulation introducing a density distribution function which is formulated as the derivative of the sigmoid function $\Phi_{s}(x)=(1+e^{-sx})^{-1}$, that is: $\phi_{s}(x)=se^{-sx}/(1+e^{-sx})^{2}$. The standard deviation of $\phi_{s}$ is $1/s$ and, as the network converges to an accurate surface representation, this parameter moves toward zero. Following NeuS’s formulation, the volumetric alpha used in Eq. 1 is replaced by the following equation to approximate the color $\hat{C}(r)$:
$\alpha_{i}=\text{max}\left(\frac{\Phi_{s}(f(p_{i}))-\ \Phi_{i}(f(p_{i+1}))}{\Phi_{s}(f(p_{i}))},0\right),$ (3)
where $f(p_{i})$ and $f(p_{i+1})$ are signed distance values at section points centered on $x_{i}$. In our formulation, we denote alpha values derived from the estimated SDF as $\alpha_{i}^{f}$.
3.2 SCILLA’s hybrid architecture
Unlike SoTA methods that singularly represent either a density or an SDF field, our hybrid method SCILLA learns the two different fields simultaneously, as illustrated in Fig 3. SCILLA expresses the scene using two functions, $\mathcal{F}_{\Theta}^{h}$ and $\mathcal{F}_{\Theta}^{c}$, encoded respectively by two MLPs. The input 3D position $x$ and viewing direction $d$ are embedded respectively via a positional and directional embedding into $\text{emb}_{x}(x)$ and $\text{emb}_{d}(d)$.
$\mathcal{F}_{\Theta}^{h}$ takes as input $\text{emb}_{x}(x)$ and outputs a density $\sigma$, a signed distance value $f(x)$, and a latent vector $h$, that is $(\sigma,f(x),h)=\mathcal{F}_{\Theta}^{h}\left(\text{emb}_{x}(x)\right).$
$\mathcal{F}_{\Theta}^{c}$ on the other hand takes as input the latent vector $h$ that is outputted by $\mathcal{F}_{\Theta}^{h}$ in addition to the embedded viewing direction $\text{emb}_{d}(d)$ and the normal vector $\overrightarrow{n}$ obtained from the gradient of the SDF, to predict the color $c=\mathcal{F}_{\Theta}^{c}\left(\text{emb}_{d}(d),h,\overrightarrow{n}\right)$.
[Uncaptioned image] Figure 3: SCILLA architecture – Our method uses two MLP functions $\mathcal{F}_{\Theta}^{h}$ and $\mathcal{F}_{\Theta}^{c}$ to output SDF values and density ($\sigma$ and $f(x)$, respectively) along with color values given an input sample. We inspire from NeRF and design $\mathcal{F}_{\Theta}^{c}$ to output the color given the lattent vector $h$ outputted by $\mathcal{F}_{\Theta}^{h}$, the viewing direction ($d$) and the normal vector $\overrightarrow{n}$ obtained from the gradient of the SDF
[Uncaptioned image] Figure 4: Sample attribution during hybrid stage – Left: we start by assigning samples with the lowest density, i.e farthest away from the real surface, to alpha values computed from the SDF (green dots) and the rest to alpha values computed from estimated density (red dots). Middle, right: as the training advances, the proportion of samples associated with SDF values increases and the signed distance field (dotted green line) can be estimated more precisely.
3.3 SCILLA’s volume rendering
3.3.1 Probabilistic density estimation.
We rely on probabilistic density estimation along the ray for a more targeted sampling strategy that samples fewer points and accelerates the overall method’s training. We sample $N$ points guided by the estimated density along the ray, where higher density indicates the presence of the surface. Given these samples, we compute using SCILLA’s architecture the alpha values and perform volume rendering to obtain the final color associated with the ray, which is compared to the ground truth value to train our model:
[…]
The strategy of associating a sample to either a volumetric or surface-based representation is discussed in the next section. The weights $w_{i}$ associated with each sample (Eq. 1) serve as a supervision signal to our probabilistic density estimator. Hence, we can refine in a self-supervised manner the prediction of the expected density along the ray. We rely on Mip-NeRF 360’s [1] proposal networks as our volumetric density estimator and use their introduced proposal loss, $\mathcal{L}_{\text{prop}}$, to refine density estimation from the weights $w_{i}$ of our main model.
3.3.2 Volumetric to surface representation.
Starting from a pure volumetric density training ($\alpha_{i}=\alpha_{i}^{v}$), we introduce a progressive transition from a volumetric to an SDF representation. Our approach is based on the premise that estimating the density $\sigma_{i}$ at a specific location requires less complex geometric understanding in contrast to estimating the signed distance to the closest point to the surface (i.e. $f(x_{i})$). By initially leveraging a pure volumetric density representation during training, we shorten the convergence time of our density estimator. Our progressive transitioning strategy can be outlined as follows:
  • Volumetric stage: all the samples along the ray express volumetric densities, and their alpha values are computed with $\sigma_{i}$.
  • Hybrid stage: we gradually assign samples along the ray to train our signed distance field. Within the same ray, a portion of sample values is represented with volumetric alpha $\alpha_{i}^{v}$, while the remainder is computed with $\alpha_{i}^{f}$.
  • Surface stage: we increase the number of samples for which alpha values are computed using $f(x)$ until we transition to a complete surface representation.
3.3.3 Hybrid stage regularization.
Composing alpha values obtained from our two different geometric representations is not straightforward as the learning dynamic of the density and the SDF are not the same. As mentioned earlier, learning a volumetric representation of the scene based only on spatial density is easier and faster to learn compared to learning an arbitrary signed distance field of a complex outdoor scene. To facilitate the convergence of our representation, we introduce two regularization methods during the hybrid stage: a probability-based sampling attribution and an SDF-gradient normalization.
Probability-based samples attribution. To assign the alpha value based on density or SDF to a sample, we initially start with a random selection for each ray. We noticed that this association strategy is not suitable: the signed distance field could not be properly initialized during the hybrid stage. To solve this problem, we introduce a probability-based sample association. We assign $\alpha_{i}^{f}$ to samples with the lowest predicted density to facilitate the learning of the signed distance field. Following this sample attribution strategy, we can estimate a coarse surface during the first iterations and refine the fine geometrical details as the training progresses, as illustrated in the toy example in Fig. 4.
SDF-gradient normalization. Volumetric representations based on density estimation tend to quickly approximate saturated alpha values ($\alpha_{i}^{v}$ being either 00 or $1$), while alpha values computed from the SDF converge slower. Because $\alpha_{i}^{v}$ and $\alpha_{i}^{f}$ are composed jointly during the hybrid stage, we noticed that surface representation compensates for this gap by predicting large SDF gradient values so that $\alpha_{i}^{f}$ aligns with $\alpha_{i}^{v}$. Indeed, from Eq. 3 a solution to saturate alpha towards either 1 or 0, is to predict $f(p_{i+1})\ll f(p_{i})$ or $f(p_{i+1})\gg f(p_{i})$, respectively. During the forward pass, $f(p_{i})$ and $f(p_{i+1})$ are not directly predicted by the model but rather computed using:
$\begin{split}f(p_{i})&=f(x_{i})+\text{Relu}(-\cos({\theta}))\times\frac{\delta_{i}}{2},\\ f(p_{i+1})&=f(x_{i})-\text{Relu}(-\cos({\theta}))\times\frac{\delta_{i}}{2},\end{split}$ (5)
with $\theta$ being the angle between the direction of $\nabla f(x_{i})$ and the ray direction $d$. Practically, we do not formally compute $\theta$ but we rather approximate its cosine with $\cos({\theta})=\nabla f(x_{i})\cdot d$, assuming both vectors are unit norm. By predicting $\left\|\nabla f(x_{i})\right\|_{2}\gg 1$, $\alpha_{i}^{f}$ can be easily saturated and follow the distribution of $\alpha_{i}^{v}$ without learning a proper signed distance field. To address this, we simply normalize the SDF gradient before the cosine computation to prevent the gradient from compensating the alpha distribution difference. It is important to notice that even if the eikonal loss used for training (see Sec. 3.4.3) is supposed to encourage the network to model a signed distance function with spatial derivative of unitary norm, we found that numerically normalizing the gradient during our hybrid stage is essential.
3.4 Optimization
3.4.1 Photometric loss.
We use the standard $L_{1}$ loss to minimize the pixel-wise color difference between the rendered image $\hat{C}$, and the ground truth image $C$, as well as DSSIM [33] loss on color patches.
3.4.2 Monocular cues.
As we target to reconstruct outdoor datasets, we add an additional sky loss to our model. Similar to StreetSurf [11], we model the sky color with an auxiliary MLP conditioned on the ray direction and use a loss to constrain the opacity of sky pixels to zero. The segmentation mask is obtained with an off-the-shelf semantic segmentation network [6].
Inspired by MonoSDF[40] and StreetSurf [11], we use off-the-shelf monocular depth estimation models to supervise the rendered normals $\hat{N}(r)$. In standard solutions [40, 11], the monocular normal loss is minimized according to the predicted normal accumulated along the ray:
$\hat{N}(r)=\sum_{i=1}^{N}T_{i}\alpha_{i}\frac{\nabla f(x_{i})}{\left\|\nabla f(x_{i})\right\|_{2}},$ (6)
$\mathcal{L}_{\hat{N}}=\left\|\hat{N}(r)-\bar{N}(r)\right\|+\left\|1-\hat{N}(r)^{\top}\bar{N}(r)\right\|.$ (7)
In our proposal, however, for more accurate and targeted supervision, we supervise only the normal associated with the closet sample $x_{N}$ to the surface estimated by transmittance saturation:
$\mathcal{L}_{\hat{N}}=\left\|\frac{\nabla f(x_{N})}{\left\|\nabla f(x_{N})\right\|_{2}}-\bar{N}(r)\right\|+\left\|1-\left(\frac{\nabla f(x_{N})}{\left\|\nabla f(x_{N})\right\|_{2}}\right)^{\top}\bar{N}(r)\right\|.$ (8)
3.4.3 Signed distance field regularization.
Similar to [34], to constrain our model to respect the eikonal equation [10] and obtain a more accurate surface reconstruction, we use the eikonal regularization defined as:
$\mathcal{L}_{\text{eik}}=\frac{1}{N}\sum_{i=1}^{N}\left(\left\|\nabla f(x_{i})\right\|_{2}-1\right)^{2}.$ (9)
According to NeuS formulation of the alpha compositing equation (Eq. 3), the network converges properly when the standard deviation $1/s$ of $\phi_{s}$ approaches zero. In complex driving scenarios and very large scenes, the parameter $s$ doesn’t always reach a satisfactory level of convergence. Unlike StreetSurf which manually fixes the parameter $s$ to increment linearly, we introduce a regularization term on $s$: $\mathcal{L}_{\text{s}}=1/(s+\epsilon)$.
4 Experiments
We evaluate the performance of SCILLA on multiple driving scenes and provide qualitative and quantitative comparisons compared to a relevant SoTA method.
4.0.1 Implementation details.
We use hash encoding to encode the positions [23] (with the same default parameters as in InstantNGP [23]), and spherical harmonics to encode the viewing directions. We use 2 layers with 64 hidden units for the MLPs $\mathcal{F}_{\Theta}^{h}$ and $\mathcal{F}_{\Theta}^{c}$. Pure volumetric stage last for the 100 first steps and the hybrid stage ends at $35\%$ of the total training time. We perform our experiments on a single high-tier GPU using Adam optimizer with a cosine learning rate decay from $10^{-2}$ to $10^{-4}$. For the parameter $s$, we apply a slightly slower weight decay ($10^{-3}$ to $10^{-5}$) to stabilize convergence. We train our model for a total of 14k iterations. Following the literature, we use Marching Cubes [19] to generate the final mesh that represents the scene. Further details on implementation can be found in the supplementary materials.
4.0.2 Datasets.
We evaluate our method on four public autonomous driving datasets: KITTI-360 [18], nuScenes [3], Waymo Open Dataset [29] and Pandaset [38]. We conduct our experiments on four scenes from each one of these datasets. Sequences were selected to have large differences between them. All sequences are static, except for Pandaset where dynamic vehicles are masked during training. Complete details about the selected sequences can be found in our supplementary.
4.0.3 Baseline.
We benchmark our method against StreetSurf [11], which is the current SoTA method for driving scenes surface reconstruction. StreetSurf relies on hierarchical space partitioning and uses 3 distinct fields to model close range, far range and sky areas. It employs both 3D and 4D hash grid positional encoding as well as occupancy grids for efficient ray sampling. We used their official open-source implementation and evaluated their method without LiDAR supervision for a fair comparison. Trajectories are not optimized, relying only on monocular cues (depths and normals) for more challenging setups. All experiments were performed on the same computational setup for fair efficiency comparisons.
4.0.4 Evaluation metrics.
To evaluate the quality of the reconstructed surfaces, we report two metrics:
[…]
  • Precision (Prec.): the percentage of LiDAR points with a distance to the mesh below $0.15$m.
4.1 Results
(a) KITTI-360 [18]
(b) Pandaset [38]
(c) Waymo [29]
(d) nuScenes [3]
Figure 5: Mean cumulative delta error for both SCILLA and StreetSurf computed across the four sequences from each dataset as detailed in table 1. Curves indicate the percentage of ground-truth points having an error distance to the closest predicted mesh triangle which is lower than a given value. Light contours represent the standard deviation for each method.
[…]
KITTI-360 [18] Pandaset [38]
Seq. 30 Seq. 31 Seq. 35 Seq. 36 Seq. 23 Seq. 37 Seq. 42 Seq. 43
[…]
StreetSurf [11] 0.14 0.50 0.09 0.71 0.10 0.67 0.10 0.66 2.33 0.10 0.47 0.40 0.36 0.30 0.28 0.31
SCILLA (ours) 0.13 0.56 0.11 0.70 0.11 0.66 0.13 0.72 0.49 0.23 0.28 0.46 0.33 0.29 0.35 0.34
Waymo Open Dataset  [29] nuScenes [3]
Seq. 10061 Seq. 13196 Seq. 14869 Seq. 102751 Seq. 0034 Seq. 0071 Seq. 0664 Seq. 0916
[…]
StreetSurf [11] 0.25 0.42 0.35 0.53 0.26 0.34 0.33 0.23 0.91 0.27 0.78 0.47 0.67 0.50 0.65 0.28
SCILLA (ours) 0.20 0.43 0.22 0.48 0.15 0.47 0.21 0.29 0.42 0.20 0.22 0.59 0.40 0.40 0.22 0.54
4.1.1 Quantitative analysis.
Table 2: Training analysis.
Iteration time GPU memory Params size
(s) (Gb) (MiB)
StreetSurf [11] 0.2 19 92.59
SCILLA (ours) 0.1 8 27
[…]
Efficiency. We report in Tab. 4.1.1 the computational performance of each method. Because SCILLA relies on a single 3D hierarchical hash grid in comparison to the two grids (3D and 4D) used in StreetSurf [11], our method is much more efficient by being designed with a third of the parameters of StreetSurf and taking half of the GPU memory. Furthermore, for a single scene, SCILLA requires half the training time compared to StreetSurf, on average.
KITTI-360 [18]
RGB Refer to caption Refer to caption Refer to caption
SCILLA (ours) Refer to caption Refer to caption Refer to caption
StreetSurf [11] Refer to caption Refer to caption Refer to caption
(a) Seq. 30 (b) Seq. 31 (c) Seq. 36
Figure 6: Qualitative experiments results on three sequences of KITTI-360. We report the RGB ground-truth image, from a different point-of-view than the RGB we compare our generated SDF mesh to StreetSurf’s mesh.
4.1.2 Qualitative analysis.
The quantitative evaluation presented in the previous section (Sec. 4.1.1) is computed from the cumulated ground-truth LiDAR which are sparse and do not cover some regions of the scene. To complete our evaluation, we present qualitative results in Figs.  6, 7, and 9. Results demonstrate that SCILLA reconstructs higher-quality surfaces compared to StreetSurf. We observe that our method recovers more accurate and realistic scene details at many regions such as buildings, cars, poles, and trees as it is highlighted by the red squares shown on each image. Our analysis supports the hypothesis that SteertSurf’s disentanglement of the close and far range based on the assumption that driving scenes are long and narrow is not adapted to the majority of driving sequences. In addition, StreetSurf’s road-surface initialization fails in sequences where the road and sidewalks are not at the same height or at non-flat roads such as downhills (see Fig. 6 – Seq. 30).
Pandaset [38]
RGB Refer to caption Refer to caption Refer to caption Refer to caption
SCILLA (ours) Refer to caption Refer to caption Refer to caption Refer to caption
StreetSurf [11] Refer to caption Refer to caption Refer to caption Refer to caption
(a) Seq. 23 (b) Seq. 37 (c) Seq. 42 (d) Seq. 43
Figure 7: Qualitative experiments results on four sequences of Pandaset. We report the masked RGB ground-truth image. In a different point-of-view than the GT we compare our generated SDF mesh to StreetSurf’s mesh.
nuScenes [3] Waymo Open Dataset [29]
RGB Refer to caption Refer to caption Refer to caption Refer to caption
SCILLA (ours) Refer to caption Refer to caption Refer to caption Refer to caption
StreetSurf [11] Refer to caption Refer to caption Refer to caption Refer to caption
(a) Seq. 71 (b) Seq. 916 (c) Seq. 10061 (d) Seq. 102751
Figure 8: Qualitative experiments on both nuScenes [3] (Left), and Waymo [29] (Right). We report the RGB ground-truth image. In a different point-of-view than the GT we compare our generated SDF mesh to StreetSurf’s mesh.
4.2 Ablation study
In this section, we ablate the effect of SCILLA’s hybrid volume rendering strategy and show quantitative and qualitative results in Tab. 3. The results demonstrate that transitioning from the volumetric to the surface representation is crucial to accurately reconstruct the road and various scene aspects such as signs and trees. In addition, we report our method with standard accumulated normal supervision (Sec. 3.4.2), and without the network regularization loss (Sec. 3.3.3).
[…]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
GT RGB (a) w/o Hybrid (b) w/o $\mathcal{L}_{\hat{N}}$ (c) w/o $\mathcal{L}_{\text{s}}$ (d) SCILLA
[…]
Prec. 0.38 0.31 0.42 0.54
5 Discussion
5.0.1 Applications.
Thanks to the quality and completeness of SCILLA’s reconstructed meshes, we can use modern MVS tools such as OpenMVS [4] to obtain a detailed and colorized representation of large urban scenes, as shown in Fig. 1. In addition, SCILLA is not restricted to driving scenarios, it can be applied to other types of scenes. We encourage the reader to refer to the supplementary materials for qualitative results of SCILLA on the Tanks&Temples dataset [16].
5.0.2 Limitations.
Although our method can achieve highly detailed surface reconstruction in driving scenarios, it fails to reconstruct fine details in very wide and open sequences. We recommend consulting the supplementary materials for the failure cases of SCILLA.
6 Conclusion
In this work we presented SCILLA, an implicit hybrid approach to efficiently reconstruct large urban scenes. Our method’s hybrid architecture models the volumetric density and signed distance in two separate fields. We introduced a novel hybrid volume rendering strategy to progressively transition from volumetric representation to a signed distance representation. The quantitative and qualitative evaluations that we performed show that SCILLA achieves more precise reconstruction than state-of-the-art methods while being significantly faster and is better suited for a large variation of driving scenarios.
Appendix A Additional implementation details
The overall loss we use to optimize SCILLA is defined as follows:
$\mathcal{L}=\mathcal{L}_{\text{rgb}}+\lambda_{1}\mathcal{L}_{\text{dssim}}+\lambda_{2}\mathcal{L}_{\hat{N}}+\lambda_{3}\mathcal{L}_{\text{eik}}+\lambda_{4}\mathcal{L}_{\text{sky}}$ (10)
We set $\lambda_{1}$, and $\lambda_{4}$ to 0.1 and 0.01, respectively. We fix $\lambda_{2}$ to 0.05 for planar classes and 0.01 for non planar classes. We set $\lambda_{3}$ to 0.01 in the first training iterations, then we adjust it to 0.1 in the last iterations.
We report in table 4 the hash grid encoding parameters from Instant-NGP [23]. We summarize the split of KITTI-360 [18] sequences used for our evaluations in table 5. We use all four cameras for KITTI-360 [18], and the three front cameras for Pandaset [38], nuScenes [3] and Waymo Open Dataset  [29].
Table 4: Hash grid encoding parameters
Parameter Value
Table size $2^{19}$
Finest resolution 2048
Coarsest resolution 16
Number of level 16
Table 5: Selected KITTI-360 sequences
Seq. Start End # frames per cam.
30 1728 1822 48
31 2890 2996 54
35 980 1092 57
36 112 166 28
Appendix B Additional results
nuScenes [3] Waymo Open Dataset [29]
RGB Refer to caption Refer to caption Refer to caption Refer to caption
SCILLA (ours) Refer to caption Refer to caption Refer to caption Refer to caption
StreetSurf [11] Refer to caption Refer to caption Refer to caption Refer to caption
(a) Seq. 34 (b) Seq. 664 (c) Seq. 13196 (d) Seq. 14869
Figure 9: Qualitative experiments on both nuScenes [3] (Left), and Waymo [29] (Right). We report the RGB ground-truth image. In a different point-of-view than the GT we compare our generated SDF mesh to StreetSurf’s mesh.
B.1 Additional qualitative results
We report in figure 9 additional qualitative results on nuScenes [3] and Waymo Open Dataset [29]. We find that SCILLA reconstructs higher-quality surfaces compared to StreetSurf and can recover many scene details (see red-boxes highlighted on the figures).
B.2 Ablation study
We ablate the effect of random sample attribution and our proposed probability-based samples attribution. The qualitative results at various training steps are presented in figure 10. The results demonstrate that SCILLA samples attribution strategy initially learns the coarse geometry of the scene during early training stages. While random sample attribution can approximate an accurate SDF representation by the end of the hybrid stage, it results in an incomplete mesh compared to the mesh generated using probability-based sample attribution.
Refer to caption Refer to caption Refer to caption
(a) Font Left RGB (b) Front RGB Front right RGB
Random Refer to caption Refer to caption Refer to caption Refer to caption
Prob-based Refer to caption Refer to caption Refer to caption Refer to caption
(a) 1.2k steps (b) 3.6k steps (c) 6k steps (d) 14k steps
Figure 10: Ablation study: we ablate the effect of random samples attribution compared to our probability-guided attribution introduced in section 3.3 of the paper. We show the rendered meshes results of the sequence 023 of Pandaset [38] at (a) 1.2k steps, (b) 3.6k steps, (c) 6k steps, and (d) 14k steps.
Appendix C Applications
C.1 Tanks And Temples dataset
In this section, we report SCILLA’s qualitative results on Tanks and Temples dataset [16]. The results reported in figure 11 demonstrate that SCILLA is not restricted to driving datasets and urban areas. SCILLA can represent large, detailed, indoor and outdoor sequences from Tanks And Temples. We report the results of SCILLA with and without the hybrid volumetric stages. Our observation suggests that the transition from the volumetric to surface representation enables to recover various scene details. In addition, SCILLA’s training time takes less than an hour, compared to other state-of-the-art indoor and outdoor landmarks reconstruction methods, such as Neuralengelo [17] which requires 10h training time.
TanksAndTemples [16]
RGB Refer to caption Refer to caption Refer to caption Refer to caption
w/o hybrid Refer to caption Refer to caption Refer to caption Refer to caption
SCILLA (ours) Refer to caption Refer to caption Refer to caption Refer to caption
(a) Auditorium (b) Temple (c) Truck (d) Meetingroom
Figure 11: Qualitative experiments on TanksAndTemples dataset [16]. We report the RGB ground-truth image. In a different point-of-view than the GT RGB, we compare the extracted mesh from the SDF from both SCILLA, and SCILLA without the hybrid stages.
C.2 Textured mesh
Due to the high-quality of SCILLA’s reconstructed surfaces, we can leverage modern Multi-View Stereo (MVS) tools like OpenMVS [4] to produce detailed and colorized representations of driving sequences. We find that SCILLA’s textured mesh is more complete and accurate compared to StreetSurf’s textured mesh.
SCILLA (ours) Refer to caption Refer to caption
StreetSurf [11] Refer to caption Refer to caption
Figure 12: We use OpenMVS [4] to colorize the outputted meshes. We compare our colored mesh to StreetSurf [11], for the sequence 102751 from Waymo Open Dataset [29].
Refer to caption Refer to caption Refer to caption Refer to caption
(a) RGB (b) Predicted sky mask (c) SCILLA (ours) (d) StreetSurf [11]
Figure 13: Failure case of SCILLA for the sequence 31 from KITTI-360 [18]. We report the ground-truth RGB image and the predicted sky mask. In a different point-of-view than the GT RGB, we compare our generated SDF mesh to StreetSurf’s mesh.
Refer to caption Refer to caption Refer to caption Refer to caption
(a) RGB (b) RGB highlighted (c) SCILLA (ours) (d) StreetSurf [11]
Figure 14: Failure case of SCILLA for the sequence 23 from Pandaset [38]. We report the ground-truth RGB image and the wide part of the sequence highlighted. In a different point-of-view than the GT RGB, we compare our generated SDF mesh to StreetSurf’s mesh.
Refer to caption Refer to caption Refer to caption Refer to caption
(a) RGB (b) Monocular normal (c) SCILLA (ours) (d) StreetSurf [11]
Figure 15: Failure case of SCILLA for the sequence 664 from nuScenes [3]. We report the ground-truth RGB image and the monocular normal. In a different point-of-view than the GT RGB we compare our generated SDF mesh to StreetSurf’s mesh.
Appendix D Limitations
Although SCILLA’s reconstructed surfaces are highly detailed and accurate, we find that our method can fail in three distinct scenarios:
  • Disentangling fine details from the sky: unlike StreetSurf [11], which models close range, far range, and sky separately, SCILLA separates only the sky from the other scene’s modeling. However, SCILLA may struggle to distinguish fine details from the sky, particularly in cases where objects are thin, as shown in figure 13.
  • Fine details in wide sequences: SCILLA sometimes fails in accurately reconstructing scene details in wide and open sequences, such as scene 23 from Pandaset reported in figure 14.
  • Inaccurate road reconstruction: for sequence 664 from nuScenes (see figure 15), SCILLA faces challenges in reconstructing the road due to inaccuracies in the monocular normal prediction from Omnidata [8]. In comparison, StreetSurf [11] demonstrates more accurate road reconstruction, attributed to its road-surface initialization.